OpenStudy (anonymous):

Find all the possible values between 0 and ten that will satisfy.... x≅1(mod 2)???? I answered 3, 5, 7 and 9.

Parth (parthkohli):

I'm just curious, but why didn't you choose 1?

OpenStudy (anonymous):

That is actually why I posted the question because I couldn't figure out why 1 satisfied the equation.

Parth (parthkohli):

I didn't understand the question :/

Parth (parthkohli):

\(\mathbb{mod}(x,y)\) Is one value missing?

OpenStudy (anonymous):

what do you mean?

Parth (parthkohli):

What is \(1(mod2)\)?

OpenStudy (anonymous):

basically I was asked to find all of the values of x that are CONGRUENT to 1(mod 2)

Parth (parthkohli):

Yes, I know, but what is \(1(\mathbf{mod} 2)\)?

OpenStudy (anonymous):

Modular arithmetic. 1(mod2), divide a number by 2 and get a remainder of 1.

Parth (parthkohli):

Do you mean that we have to find the values of x where \(mod(x,2) = 1\)? I think that's your question.

Parth (parthkohli):

Yes, I know it. And indeed, that's you question.

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow x \epsilon 1,3,5,7,9 }\)

OpenStudy (anonymous):

what notation are you using?

Parth (parthkohli):

\(\large \color{Black}{\Rightarrow x \text{ } \epsilon \text{ } \mathfrak{ odd } }\)

OpenStudy (anonymous):

How did you get 1?

Parth (parthkohli):

Yeah I'm sorry, I couldn't understand the notation.

Parth (parthkohli):

0 divided by 2 leaves remainder 1 and quotient 0.

OpenStudy (anonymous):

How did you solve this problem may I ask?

Parth (parthkohli):

And you should've used \(\equiv\) instead of \(\cong\). Hope this helped.

Parth (parthkohli):

Well you check all the values that give the remainder as 1.

OpenStudy (anonymous):

0 divided by 2 leaves remainder 1 and quotient 0!!!, what type of maths of this?

Parth (parthkohli):

Lol it's the same thing.

OpenStudy (anonymous):

lol that has really confused me.

Parth (parthkohli):

You have to check the numbers that leave the remainder as 1, and 0/2 leaves the remainder as 1. agree?

OpenStudy (anonymous):

my calculator does not agree no.

Parth (parthkohli):

Oopsie. I got confused lol.

Parth (parthkohli):

I meant 1/2

OpenStudy (anonymous):

I am quite confused about how 1 can be a solution.

Parth (parthkohli):

All odd numbers are solutions.

OpenStudy (anonymous):

what about about the whole remainder of 1 business?

Parth (parthkohli):

That is the question.

Parth (parthkohli):

x = 1(mod 2) means that when x is divided by 2, it leaves the remainder of 1.

OpenStudy (anonymous):

and 1/2=1/2, so I cannot understand now 1 can be a solution.

Parth (parthkohli):

All the numbers that can be expressed as \(2\mathbb{Z} + 1\) are solutions here.

Parth (parthkohli):

\(\mathbb{Z}\) can be 0 here. So, 2(0) + 1 = 1

Parth (parthkohli):

Divide 1 and 2, it'll leave a remainder of 1.

OpenStudy (anonymous):

could how show me please?

Parth (parthkohli):

http://www.wolframalpha.com/input/?i=Remainder+of+1+over+2 See this.

Parth (parthkohli):